OpenStudy (anonymous):

Can the expression s sqrt(3) + sqrt(27s^5) + 9 sqrt(3s) be simplified into an expression with fewer terms? Explain and simplify, if possible.

OpenStudy (mertsj):

\[\sqrt{3}+\sqrt{27s^5}+9\sqrt{3s}=\sqrt{3}+\sqrt{9(3)s^4s}+9\sqrt{3s}=\sqrt{3}+3s^2\sqrt{3s}+9\sqrt{3s}=\]

OpenStudy (anonymous):

Thanks so much!! Your alot of help!!

OpenStudy (mertsj):

\[\sqrt{3}+(3s^2+9)(\sqrt{3s})\]

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

How did you get that answer? Just asking :)

OpenStudy (mertsj):

Factor out the common factor sqrt3s

OpenStudy (anonymous):

Oh okay thanks!

OpenStudy (mertsj):

yw